OpenStudy (calculusxy):

Medal for a geometry question!!! Use the given information to write an equation and solve the problem. 1. A supplement of an angle is six times as large as a complement of the angle. Find the measures of the angle, its supplement, and its complement.

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let x = the unknown angle

jimthompson5910 (jim_thompson5910):

the `supplement of an angle ` means we have some other angle, say y, such that x+y = 180 so y = 180-x

jimthompson5910 (jim_thompson5910):

agreed so far?

OpenStudy (calculusxy):

Would it be like complementary = 90 - x supplementary = 180 - x 90 - x = 6(90 - x) 90 - x = 540 - 6x

OpenStudy (calculusxy):

Sorry I meant 180 - x = 6(90 - x) 180 - x = 540 - 6x

jimthompson5910 (jim_thompson5910):

yes the starting equation would be 180 - x = 6(90 - x)

OpenStudy (calculusxy):

And then -360 - x = -6x -360 = -5x 72 = x

OpenStudy (calculusxy):

Angle = 72 degrees Complementary = 18 degrees Supplementary = 108 degrees

OpenStudy (calculusxy):

Are those correct?

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have the correct answer

OpenStudy (calculusxy):

Thanks for the help :)

jimthompson5910 (jim_thompson5910):

You're welcome